Solve the following system of linear equations. (If the system has no solution, enter NONE for each answer. If the system is dependent, enter INFINITELY MANY for each answer.)
x+5y=2
y=x+4
x = ?
y = ?
Pls show me the steps on how to get the answer. thanks

Answer by shree840(260)

step 1 put both the eq in the same form
x+5y=2
x-y=-4
------- we have several ways
1. By substitution
x=2-5y from eq 1 and substitute in eq 2
2-5y-y=-4
-6y=-6
y=1
x=2-5y=2-5=-3